OpenStudy (anonymous):

the length of a rectangular boat storage shed must be 4 meters more than the width and the perimeter must be at least 120 meters. what is the range of values for the width? exactly 28 meters? at least 28 meters? at most 28 meters?

OpenStudy (anonymous):

@MRamosville

OpenStudy (anonymous):

If your length is x, than your width must be (x - 4). x+x + (x -4) + (x - 4) = 120 4x - 8 = 120 4x +8 = + 8 4x = 128 4x/4 = 128/8 x = 32 Your length is 32, which means your width must at least be 28. It could be more, but it can't be less.

OpenStudy (anonymous):

Your length and your width could both be bigger, but the width is at least 28.